# Detecting and Mitigating System-Level Anomalies of Vision-Based Controllers

### v3

@@ Abstract
Autonomous systems, such as self-driving cars and drones, have made significant strides in recent years by leveraging visual inputs and machine learning for decision-making and control. Despite their impressive performance, these vision-based controllers can make erroneous predictions when faced with novel or out-of-distribution inputs. Such errors can cascade to catastrophic system failures and compromise system safety. In this work, we introduce a run-time anomaly monitor to detect and mitigate such closed-loop, system-level failures. Specifically, we leverage a reachability-based framework to stress-test the vision-based controller offline and mine its system-level failures. This data is then used to train a classifier that is leveraged online to flag inputs that might cause system breakdowns. The anomaly detector highlights issues that transcend individual modules and pertain to the safety of the overall system. We also design a fallback controller that robustly handles these detected anomalies to preserve system safety. We validate the proposed approach on an autonomous aircraft taxiing system that uses a vision-based controller for taxiing. Our results show the efficacy of the proposed approach in identifying and handling system-level anomalies, outperforming methods such as prediction error-based detection, and ensembling, thereby enhancing the overall safety and robustness of autonomous systems.
@@ I Introduction
Anomaly detection (AD) methods for learning components typically fall under two categories: distributional shift methods and functional uncertainty methods [3]. The former aims to detect and mitigate distribution shifts between training and the test times. These methods include approaches that artificially inject noise in the ground truth data (e.g., images) to force a distribution shift and create anomalous data [4]. Another line of research in this direction aims to develop learning algorithms that are distributionally robust, optimizing the worst-case performance within a pre-specified envelope of distributional shifts to guarantee out-of-distribution performance [5, 6] or performing domain randomization during training [7]. Since detecting distribution shifts can be challenging in general, especially when the test distribution is unknown a priori, the functional uncertainty methods instead detect the inputs that are either dissimilar to the training data [8, 9, 10, 11, 12] or lead to erroneous or low-confidence predictions [13, 14, 15, 16, 17, 18]. The above methods predominately detect anomalous inputs at the component level; however, such component level monitoring (e.g., detecting image classification errors) can be insufficient to prevent system-level faults. Seemingly minor errors in the individual modules can cascade into catastrophic effects at a system level. Hence, a system-level view of such problems is often encouraged and is the main motivation behind this work.
@@ II Problem Setup
Consider a robot with dynamics, $\dot{\mathbf{x}}=f(\mathbf{x},u)$ where, state $\mathbf{x}\in\mathbb{R}^{n}$ and control $u\in\mathbb{U}$ (a compact set). The robot possesses a sensor $S$, that allows it to perceive visual inputs from its surroundings at any given state $\mathbf{x}$. We can express the mapping between the state to the input as, $I=S(\mathbf{x})$. $I$ could be an RGB image, a pointcloud, etc. Additionally, we have a vision-based controller, $\uppi$, that maps $I$ to the control $u$, and defined as, $u\coloneqq\uppi(I)$. Note that $\uppi$ can be an end-to-end policy or it can consists of several sub-modules some of which may not be data-driven.
@@
Running example (TaxiNet). We introduce the autonomous aircraft taxiing problem [20] as a running example to illustrate the key aspects of our framework. Here, the robot is a Cessna 208B Grand Caravan aricraft modeled as a three-dimensional non-linear system with dynamics:
@@ V Case Study: Autonomous Aircraft Taxiing
To compute the BRTs, we use the Level Set Toolbox (LST)[24] that solves the HJB-VI in (6) numerically over a state-space grid for a time horizon of 8s. Following [19], we use a uniform $101\times 101\times 101$ grid over $p_{x}\in[-X,X]m$, $p_{y}\in[100,250]m$ and $\theta\in[-28\degree,28\degree]$ (the value of $X$ depends on the runway, e.g., for KMWH, X = 11). The LST requires the control input $u$ at each of the grid point to compute the BRT, which is obtained by rendering the image using X-Plane simulator at that state and querying the controller in (3). This consists of passing the rendered image through the TaxiNet CNN, followed by a P-controller. We refer the readers to [19] for more details on the BRT computation.
Fig. 3 (left) shows a slice of the computed BRT for the KMWH runway at 9am in clear conditions for $p_{y}=100m$. The gray area represent the set of starting states of the aircraft from which it will eventually leave the runway under TaxiNet, whereas white represents the safe area. We next randomly sample 20K states for each of the 12 conditions and render the images for these states, leading to an overall training dataset of 240K images. If an image is generated from a state present in the BRT, we label it as an anomaly. Otherwise, we label it as safe.
@@
Anomaly Detector. We next train a binary classifier on the collected dataset that takes a 224x224x3 input image and returns the probability of the image being anomalous. We use a pre-trained EfficientNet-B0 Model [25] as our backbone, replacing the last layer with a fully connected layer that feeds into a binary softmax output layer. We trained the classifier using cross-entropy loss and Adam optimizer for 20 epochs with a fixed learning rate of $3e^{-4}$ which took around 1 hour on a Nvidia RTX 3090 GPU.
Evaluation. To test the generalization capabilities of the learned AD, we test it on an unseen time of the day (5 PM) for the three airports present in the training set, as well as on two unseen KSFO and KEWR airports, across both cloud conditions (clear and overcast). Specifically, we measure the recall and accuracy of the learned AD. A higher recall value tells us that our AD can reliably detect true positives, i.e., the AD returns a conservative prediction if it is unsure of an input. Such a behavior err on caution to produce a safety-first detector. To compute these metrics, we also obtain the BRT of the system for the test environments; however, note that during testing, our network has access to only the sampled image without any privileged information regarding the system, surroundings, or the BRTs.
@@
We summarize the performance of the anomaly detector in Table I (Here (C) = cloudy, (O) = overcast). As evident from the results, the learned AD is consistently able to detect system-level anomalies in new environments. We illustrate some of the representative images that were classified as anomalous in Fig. 4. The AD is able to learn the safe limits of the runway and that if the system is close to the runway boundary. TaxiNet may not be able to correctly estimate aircraft’s state and it might fail under the vision-based controller.
@@
This may not be particularly surprising as most states near the runway boundary (white lines in Fig. 4 (a-b)) are unsafe in the training dataset (the region near the runway boundary is contained in the BRT in Fig. 3), potentially allowing the AD to learn that such images often lead to a failure. However, what’s interesting is that the AD learns that a similar image will not cause a system failure during night time because of the runway lights, which when are lit, help the TaxiNet to estimate the aircraft position more accurately, avoiding a failure. Finally, we noticed that several images classified as anomalies contained the aircraft runway markings (Fig. 4 (d-f)). Such an emergent understanding of semantic failure mode is quite impressive since in the prior work [19], the authors had to manually analyze the failures to recognize these patterns as anomalies. Our AD justifies the use of a learned classifier by practically automating the detection process without any inclusion of heuristics or manual intervention.
@@ V-2 Ensembled Predictions
Fallback Mechanism. Equipped with a capable AD, we designed a simple fallback mechanism to ensure our aircraft’s safety under anomalous inputs. If at any point in its trajectory, the aircraft observes an image $I$ that is classified as an anomaly (i.e., $\sigma(I)=1$), the linear velocity of the plane ($v$ in Eqn. (2)) is reduced by $0.01m/s$. Intuitively, this results in slowing the aircraft every time it encounters an anomaly, ultimately coming to a complete stop if it continues to encounter anomalies. As soon as the aircraft detects an image that is non-anomalous, the default TaxiNet controller takes over. Fig. 7 shows trajectories of the aircraft under the resultant controller (red trajectories). The nominal trajectories without the fallback controller are shown in dashed black. In the first case (Fig. 7 (a)), the anomaly is triggered due to the aircraft approaching too close to the runway boundary, while in the second case (Fig. 7 (b)) the anomaly is triggered due to semantic failure of the runway markings. In both cases, the TaxiNet controller leads the system off the runway. On the other hand, the fallback controller decreases the aircraft velocity whenever an anomaly is triggered to ensure system safety. This can also be seen from the velocity variation along the red trajectories. To further illustrate the advantage of the fallback controller, we compute the system BRT under the Taxinet controller and the fallback controller pipeline for one of the test cases on KMWH runway (Fig.7(c)). The BRT (i.e., the set of failure states) is much smaller under the safety pipeline (the BRT volume decreases from 28.23% to 16.86%), showing that the proposed mechanism significantly reduces the number of closed-loop system failures.
@@
Incremental Training. A complementary approach to employing an AD can be to perform targeted incremental training of the TaxiNet model on the collected anomalous data. Essentially, these labeled anomalies represent scenarios where TaxiNet fails to perform optimally. Therefore, by conducting specialized training on these failure instances, we aim to fortify TaxiNet’s robustness in handling such cases. We present some preliminary result of incremental training of TaxiNet in Fig. 9. We plan on exploring this method further in future work.
@@ VI Discussion and Future Work
In this work, we present an approach aimed at identifying and mitigating system-level anomalies of autonomous systems that rely on vision-based controllers for decision-making. By leveraging insights from reachability analysis, our approach learns an anomaly detector that effectively tackles concerns related to system-level safety during runtime. The learned anomaly detector is combined with a fallback controller to reduce the potential for catastrophic system failures significantly.
In the future, we plan to test our approach over a wide range of real-world scenarios and focus on more intricate fallback mechanisms that can efficiently work for different system-level failures. In addition, to further enhance the performance of the AD, one can take advantage of temporal information, such as the visual history, instead of only the current image. Finally, instead of relying on a classifier, other approaches could explore unsupervised methods like clustering via support vector machines or k-means applied to our annotated dataset.

### v4

@@ Abstract
Autonomous systems, such as self-driving cars and drones, have made significant strides in recent years by leveraging visual inputs and machine learning for decision-making and control. Despite their impressive performance, these vision-based controllers can make erroneous predictions when faced with novel or out-of-distribution inputs. Such errors can cascade to catastrophic system failures and compromise system safety. In this work, we introduce a run-time anomaly monitor to detect and mitigate such closed-loop, system-level failures. Specifically, we leverage a reachability-based framework to stress-test the vision-based controller offline and mine its system-level failures. This data is then used to train a classifier that is leveraged online to flag inputs that might cause system breakdowns. The anomaly detector highlights issues that transcend individual modules and pertain to the safety of the overall system. We also design a fallback controller that robustly handles these detected anomalies to preserve system safety. We validate the proposed approach on an autonomous aircraft taxiing system that uses a vision-based controller for taxiing. Our results show the efficacy of the proposed approach in identifying and handling system-level anomalies, outperforming methods such as prediction error-based detection and ensembling, thereby enhancing the overall safety and robustness of autonomous systems.
@@ I Introduction
Anomaly detection (AD) methods for learning components typically fall under two categories: distributional shift methods and functional uncertainty methods [3]. The former aims to detect and mitigate distribution shifts between training and test times. These methods include approaches that artificially inject noise in the ground truth data (e.g., images) to force a distribution shift and create anomalous data [4]. Another line of research in this direction aims to develop learning algorithms that are distributionally robust, optimizing the worst-case performance within a pre-specified envelope of distributional shifts to guarantee out-of-distribution performance [5, 6] or performing domain randomization during training [7]. Since detecting distribution shifts can be challenging in general, especially when the test distribution is unknown a priori, the functional uncertainty methods instead detect the inputs that are either dissimilar to the training data [8, 9, 10, 11, 12] or lead to erroneous or low-confidence predictions [13, 14, 15, 16, 17, 18]. The above methods predominately detect anomalous inputs at the component level; however, such component level monitoring (e.g., detecting image classification errors) can be insufficient to prevent system-level faults. Seemingly minor errors in the individual modules can cascade into catastrophic effects at a system level. Hence, a system-level view of such problems is often encouraged and is the main motivation behind this work.
@@ II Problem Setup
Consider a robot with dynamics, $\dot{\mathbf{x}}=f(\mathbf{x},u)$ where, state $\mathbf{x}\in\mathbb{R}^{n}$ and control $u\in\mathbb{U}$ (a compact set). The robot possesses a sensor $S$, that allows it to perceive visual inputs from its surroundings at any given state $\mathbf{x}$. We can express the mapping between the state to the input as $I=S(\mathbf{x})$. $I$ could be an RGB image, a pointcloud, etc. Additionally, we have a vision-based controller, $\uppi$, that maps $I$ to the control $u$, and defined as, $u\coloneqq\uppi(I)$. Note that $\uppi$ can be an end-to-end policy, or it can consist of several sub-modules, some of which may not be data-driven.
@@
Running example (TaxiNet). We introduce the autonomous aircraft taxiing problem [20] as a running example to illustrate the key aspects of our framework. Here, the robot is a Cessna 208B Grand Caravan aircraft modeled as a three-dimensional non-linear system with dynamics:
@@ V Case Study: Autonomous Aircraft Taxiing
To compute the BRTs, we use the Level Set Toolbox (LST)[24] that solves the HJB-VI in (6) numerically over a state-space grid for a time horizon of 8s. Following [19], we use a uniform $101\times 101\times 101$ grid over $p_{x}\in[-X,X]m$, $p_{y}\in[100,250]m$ and $\theta\in[-28\degree,28\degree]$ (the value of $X$ depends on the runway, e.g., for KMWH, X = 11). The LST requires the control input $u$ at each of the grid points to compute the BRT, which is obtained by rendering the image using X-Plane simulator at that state and querying the controller in (3). This consists of passing the rendered image through the TaxiNet CNN, followed by a P-controller. We refer the readers to [19] for more details on the BRT computation.
Fig. 3 (left) shows a slice of the computed BRT for the KMWH runway at 9am in clear conditions for $p_{y}=100m$. The gray area represents the set of starting states of the aircraft from which it will eventually leave the runway under TaxiNet, whereas the white represents the safe area. We next randomly sample 20K states for each of the 12 conditions and render the images for these states, leading to an overall training dataset of 240K images. If an image is generated from a state present in the BRT, we label it as an anomaly. Otherwise, we label it as safe.
@@
Anomaly Detector. We next train a binary classifier on the collected dataset that takes a 224x224x3 input image and returns the probability of the image being anomalous. We use a pre-trained EfficientNet-B0 Model [25] as our backbone, replacing the last layer with a fully connected layer that feeds into a binary softmax output layer. We trained the classifier using cross-entropy loss and Adam optimizer for 20 epochs with a fixed learning rate of $3e^{-4}$, which took around 1 hour on a Nvidia RTX 3090 GPU.
Evaluation. To test the generalization capabilities of the learned AD, we test it at an unseen time of the day (5 PM) for the three airports present in the training set, as well as on two unseen KSFO and KEWR airports, across both cloud conditions (clear and overcast). Specifically, we measure the recall and accuracy of the learned AD. A higher recall value tells us that our AD can reliably detect true positives, i.e., the AD returns a conservative prediction if it is unsure of an input. Such a behavior err on caution to produce a safety-first detector. To compute these metrics, we also obtain the BRT of the system for the test environments; however, note that during testing, our network has access to only the sampled image without any privileged information regarding the system, surroundings, or the BRTs.
@@
We summarize the performance of the anomaly detector in Table LABEL:table:table2 (Here (C) = cloudy, (O) = overcast). As evident from the results, the learned AD is consistently able to detect system-level anomalies in new environments. We illustrate some of the representative images that were classified as anomalous in Fig. 4. The AD is able to learn the safe limits of the runway and that if the system is close to the runway boundary. TaxiNet may not be able to correctly estimate aircraft’s state and it might fail under the vision-based controller.
@@
This may not be particularly surprising as most states near the runway boundary (white lines in Fig. 4 (a-b)) are unsafe in the training dataset (the region near the runway boundary is contained in the BRT in Fig. 3), potentially allowing the AD to learn that such images often lead to a failure. However, what’s interesting is that the AD learns that a similar image will not cause a system failure during night time because of the runway lights, which, when are lit, help the TaxiNet to estimate the aircraft position more accurately, avoiding a failure. Finally, we noticed that several images classified as anomalies contained the aircraft runway markings (Fig. 4 (d-f)). Such an emergent understanding of semantic failure mode is quite impressive since in the prior work [19], the authors had to manually analyze the failures to recognize these patterns as anomalies. Our AD justifies the use of a learned classifier by practically automating the detection process without any inclusion of heuristics or manual intervention.
@@ V-2 Ensembled Predictions
Fallback Mechanism. Equipped with a capable AD, we designed a simple fallback mechanism to ensure our aircraft’s safety under anomalous inputs. If at any point in its trajectory, the aircraft observes an image $I$ that is classified as an anomaly (i.e., $\sigma(I)=1$), the linear velocity of the plane ($v$ in Eqn. (2)) is reduced by $0.01m/s$. Intuitively, this results in slowing the aircraft every time it encounters an anomaly, ultimately coming to a complete stop if it continues to encounter anomalies. As soon as the aircraft detects an image that is non-anomalous, the default TaxiNet controller takes over. Fig. 7 shows trajectories of the aircraft under the resultant controller (red trajectories). The nominal trajectories without the fallback controller are shown in dashed black. In the first case (Fig. 7 (a)), the anomaly is triggered due to the aircraft approaching too close to the runway boundary, while in the second case (Fig. 7 (b)) the anomaly is triggered due to semantic failure of the runway markings. In both cases, the TaxiNet controller leads the system off the runway. On the other hand, the fallback controller decreases the aircraft velocity whenever an anomaly is triggered to ensure system safety. This can also be seen from the velocity variation along the red trajectories. To further illustrate the advantage of the fallback controller, we compute the system BRT under the Taxinet controller and the fallback controller pipeline for one of the test cases on KMWH runway (Fig.7(c)). The BRT (i.e., the set of failure states) is much smaller under the safety pipeline (the BRT volume decreases from 25.75% to 18.24%), showing that the proposed mechanism significantly reduces the number of closed-loop system failures.
@@
Incremental Training. A complementary approach to employing an AD can be to perform targeted incremental training of the TaxiNet model on the collected anomalous data. Essentially, these labeled anomalies represent scenarios where TaxiNet fails to perform optimally. Therefore, by conducting specialized training on these failure instances, we aim to fortify TaxiNet’s robustness in handling such cases. We present some preliminary results of incremental training of TaxiNet in Fig. 9. We plan on exploring this method further in future work.
@@ VI Discussion and Future Work
In this work, we present an approach aimed at identifying and mitigating system-level anomalies of autonomous systems that rely on vision-based controllers for decision-making. By leveraging insights from reachability analysis, our approach learns an anomaly detector that effectively tackles concerns related to system-level safety during runtime. The learned anomaly detector is combined with a fallback controller to reduce the potential for catastrophic system failures significantly. In the future, we plan to test our approach over a wide range of real-world scenarios and focus on more intricate fallback mechanisms that can efficiently work for different system-level failures. In addition, to further enhance the performance of the AD, one can take advantage of temporal information, such as the visual history, instead of only the current image. Finally, instead of relying on a classifier, other approaches could explore unsupervised methods like clustering via support vector machines or k-means applied to our annotated dataset.
We introduce a framework for automatically discovering the closed-loop failures of vision-based controllers. Our work combines simulation-based approaches with HJ reachability analysis to systematically and tractably find these failures. We demonstrate the efficacy of our approach on two distinct applications – a 5-dimensional wheeled robot and a 3-dimensional aircraft using RGB-image-based neural network controllers. Our work suggests a number of interesting future directions. First, the grid-based approach to computing reachable sets suffers from the curse of dimensionality. In the future, it would be interesting to consider alternative sampling-based reachability methods, such as DeepReach [26], that are shown to scale well to high-dimensional systems. We would also like to explore using the obtained failures in improving the performance of the vision-based controller, e.g., through incremental training on the obtained scenarios. Finally, even though the failure discovery is automatic, the analysis of the failures is primarily performed manually in the current work. Automating the failure analysis, e.g., through generative and clustering methods, will be a promising future direction.